a chemist is studying the following equilibrium, which has a given equilibrium constant at a certain temperature, N2+2H2O->2H2+2NO Kp=1.x10^-8 he then fills a reaction vessel at this temperature with 13. atm of nitrogen gas and 10. atm of water vapor.
Can you predict the equilibrium pressure of NO? yes/no
If yes, what is the equilibrium pressure of NO?
Answer #1

N2 + 2H2O <=> 2H2 + 2NO

13 10 0. 0

13-x . 10-2x . 2x. 2x

Kp = (H2)^2*(NO)^2 / (N2)*(H2O)^2

1.0*10^-8 = (2x)*(2x) / (13-x)*(10-2x)

x = 5.7*10^-4

NO = 2x

= 2*(5.7*10^-4)

= 1.14*10^-3 atm

If you have any query please comment

If you satisfied with the solution please like it thankxx